OpenStudy (anonymous):

Simplify 3^4(log_3(2x))

OpenStudy (turingtest):

use exponent laws\[x^{ab}=(x^a)^b=(x^b)^a\]

OpenStudy (anonymous):

is this \[3^{4\log_3(2x)}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then start with the fact that \[4\log_3(2x)=\log_3((2x)^4)=\log_3(16x^4)\]

OpenStudy (anonymous):

Okay.. What next?

OpenStudy (anonymous):

\[b^{\log_b(z)}=z\]

OpenStudy (anonymous):

so \[3^{\log_3(16x^4)}=16x^4\]

OpenStudy (anonymous):

Okay, thank you!